OpenStudy (anonymous):

three algebra 1 students are comparing how fast their social media posts have spread. their results are shown in the following table .

OpenStudy (anonymous):

Student Amber Ben Carter Description Amber shared her photo with 3 people. They continued to share it, so the number of shares increases every day as shown by the function. Ben shared his post with 2 friends. Each of those friends shares with 3 more every day, so the number of shares triples every day. Carter shared his post with 10 friends, who each share with only 2 people each day.

OpenStudy (anonymous):

1. write an exponential function to represent the spread of bens social media post

OpenStudy (anonymous):

2. write and exponential function to represent the spread of charters social media post 3. graph each function using at least 3 points for each line. all graphs should be placed together on the same coordinate plane. 4. using the functions for each student, predict how many shares each students post will have received on day 3 and then on day 10 5. If Amber decides to mail copies of her photo to the 45 residents of her grandmothers assisted living facility, the new function representing her photo share is f(x)=3(4)^x+45 . how does this graph compare with the original graph of ambers photo share? 6. based on your results, which students post travels the fastest? how is this shown in the equation form of the functions? 7. if you had to choose , would you prefer a post with fewer friends initially but more shares, like Amber, or more friends initially but few shares?

OpenStudy (anonymous):

can someone please help I need to get this done

OpenStudy (anonymous):

@reddisja ??

OpenStudy (anonymous):

@easyaspi314

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@shirley27

OpenStudy (anonymous):

did you ever get the answer?

OpenStudy (anonymous):

Are you in FLVS?

OpenStudy (anonymous):

I am in flvs and no I never got an answer /:

OpenStudy (anonymous):

@enchanted_bubbles @EmilyF03

OpenStudy (jadatamiamaraj):

1. Write an exponential function to represent the spread of Ben's social media post. Ben=f(x)=2∗(3)x 2. Write an exponential function to represent the spread of Carter's social media post. Carter=f(x)=10∗(2)^x 3. Graph each function using at least 3 points for each line. All graphs should be placed together on the same coordinate plane, so be sure to label each line. You may graph your equation by hand on a piece of paper and scan your work or you may use graphing technology. 4. Using the functions for each student, predict how many shares each student's post will have received on Day 3 and then on Day 10. Justify your answers. Amber=f(x)=3∗(4)^x Ben=f(x)=2∗(3)x Carter=f(x)=10∗(2)^x On 3rd day : Amber=f(3)=3∗(4)^3 = 192 Ben=f(x)=2∗(3)^3 = 54 Carter=f(x)=10∗(2)^3 = 80 On 10th day : Amber=f(10)=3∗(4)^10 = 3145728 Ben=f(10)=2∗(3)^10 = 118098 Carter=f(10)=10∗(2)^10 = 10240 Amber's shares are increasing by 4 times each day where as Carter's shares are only increasing by 2 times each day. So Amber's shares are expected to be more on day 10 compared to Carter's 5. If Amber decides to mail copies of her photo to the 45 residents of her grandmother's assisted living facility, the new function representing her photo shares is f(x) = 3(4)x + 45. How does this graph compare with the original graph of Amber's photo share? This graph compares to the original graph of Amber’s photo share is both of them are exponential growths. The new function goes 45 units up from the original function which was f(x)=3(4)^x 6. Based on your results, which students' post travels the fastest? How is this shown in the equation form of the functions? Notice that Amber has started with just 3 shares and Carter started with 10 shares. Still Carter's shares are very low compared to Amber on day 10. Amber's shares are increasing by 4 times each day where as Carter's shares are only increasing by 2 times each day. So Amber's shares are expected to be more on day 10 compared to Carter's 7. If you had to choose, would you prefer a post with fewer friends initially but more shares, like Amber, or more friends initially but fewer shares? Justify your answer with your calculations from previous questions I would prefer a post with few friends initially but more shares. Having more shares implies that the base of the exponential function is greater, and this will cause posts to increase faster. We can see this in results I found; Amber's posts are greater than Ben and Carter's posts after 3 days and after 10 days.

OpenStudy (jadatamiamaraj):

https://www.desmos.com/calculator/3t1miijac8 ^^^ thats the graph for number 3